OpenStudy (anonymous):

integral of (1+v^2)^-1/2 dv

hartnn (hartnn):

heard of hyperbolic functions ?

OpenStudy (anonymous):

\[\int\limits (1/\sqrt{1+v^2}) dv\]

hartnn (hartnn):

yeah, got that. put v= sinh x

hartnn (hartnn):

dv=?

OpenStudy (anonymous):

-cosh x +c

hartnn (hartnn):

for dv ? for dv u don't need +c and dv = cosh x dx

hartnn (hartnn):

and 1+ sinh^2 x =?

OpenStudy (anonymous):

cosh^2 x

hartnn (hartnn):

dv/sqrt(1+v^2) = cosh x dx /sqrt(1+ sinh^2 x) =cosh x dx/sqrt(cosh^2 x) = cosh x dx/cosh x = dx so what is integral of dx ?

hartnn (hartnn):

got that ^ ?

OpenStudy (anonymous):

still chewing

hartnn (hartnn):

which step u have doubt with ?

OpenStudy (anonymous):

Integral of dx is just x + c

hartnn (hartnn):

yes. so we had sinh x =v what is x from here ?

OpenStudy (anonymous):

x = sinh^-1 v

hartnn (hartnn):

so your integral equals sinh^-1 v+c got it ?

OpenStudy (anonymous):

Ok. I am looking in a text book and their answer looks different. It actually stems from a larger problem. Here is the whole thing: Solve the differential equation: \[dv/\sqrt{1+v^2}=dx/x\]

OpenStudy (anonymous):

This is after rearranging to separate variables...

hartnn (hartnn):

and u know how to integrate dx/x right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ln(x) + c

hartnn (hartnn):

so u have sinh^-1 v = ln x+c

OpenStudy (anonymous):

But their depiction of "integrating both sides" looks like this:\[\ln(v+\sqrt{1+v^2}) = \ln(x) + \ln(c)\]

hartnn (hartnn):

yes, sinh^-1 v can also be written like that, ln(v+sqrt(1+v^2))

hartnn (hartnn):

both the forms are identical

OpenStudy (anonymous):

Ok, so that is a trig property that I should know (or at least can look up)?

hartnn (hartnn):

yes, u can derive that, if u want using the definition of sinh x

OpenStudy (anonymous):

Ok, got it. thank you for your help!

hartnn (hartnn):

welcome ^_^